### v1

# A1: Asynchronous Test-Time Scaling via Conformal Prediction

Abstract
Large language models (LLMs) benefit from test-time scaling, but existing methods face significant challenges, including severe synchronization overhead, memory bottlenecks, and latency, especially during speculative decoding with long reasoning chains. We introduce A1 (Asynchronous Test-Time Scaling), a statistically guaranteed adaptive inference framework that addresses these challenges. A1 refines arithmetic intensity to identify synchronization as the dominant bottleneck, proposes an online calibration strategy to enable asynchronous inference, and designs a three-stage rejection sampling pipeline that supports both sequential and parallel scaling. Through experiments on the MATH, AMC23, AIME24, and AIME25 datasets, across various draft–target model families, we demonstrate that A1 achieves a remarkable 56.7x speedup in test-time scaling and a 4.14x improvement in throughput, all while maintaining accurate rejection-rate control, reducing latency and memory overhead, and no accuracy loss compared to using target model scaling alone. These results position A1 as an efficient and principled solution for scalable LLM inference. We have released the code at https://github.com/menik1126/asynchronous-test-time-scaling.
1 Introduction
Recent advances in large language models (LLMs) drive interest in thinking models—systems that go beyond next-token prediction to emulate human-like reasoning processes (Guo et al., 2025; Muennighoff et al., 2025; McCoy et al., 2024; Shao et al., 2024). These models excel at leveraging complex reasoning chains, especially in test-time scaling settings (Snell et al., 2024; Li et al., 2025; Muennighoff et al., 2025; Zeng et al., 2025).
Test-time scaling (Chen et al., 2025; Muennighoff et al., 2025; Guo et al., 2025) constitutes an entirely new paradigm that enhances the model’s reasoning capabilities by allocating additional computational resources during the inference stage. Typically, test-time scaling can be categorized into two approaches: sequential scaling (Muennighoff et al., 2025; Guo et al., 2025) and parallel scaling (Chen et al., 2025). However, despite its potential, the challenge of efficiently managing increasing input size or complexity during inference remains a critical limitation, hindering the achievement of high-performance deployment.
Figure 1: Memory Overhead vs. Sampling Sizes (QwQ 32B, Token Budget 500)
Parallel scaling, due to its high efficiency in parallel expansion, emerges as a promising paradigm. By parallelly scaling the number of samples, Chen et al. (2025) partially alleviates the high inference latency and memory bottlenecks caused by scaling the token budget reasoning paths, while improving the efficiency of token sampling. Although some methods that adopt confidence-based early stopping of reasoning chains (Huang et al., 2025; Wan et al., 2024) improve parallel sampling efficiency, problems still remain in memory efficiency and high inference latency. Another potential issue is that early stopping prunes away potentially correct reasoning paths and reduces the diversity of the output space.
Speculative decoding (Li et al., 2024a; Leviathan et al., 2023; Kim et al., 2023; Pan et al., 2025b; Yang et al., 2025) represents a promising approach for accelerating test-time scaling. In this framework, a lightweight draft model generates tokens, which are subsequently validated by a target model. This dual-phase approach not only speeds up inference by offloading most of the generation process to the draft model, but also ensures that the final outputs retain high fidelity, thereby achieving a favorable balance between efficiency and accuracy.
However, when speculative decoding (Pan et al., 2025b; Yang et al., 2025) meets test-time scaling, the decoding process faces two key challenges. The first is the prefill stage memory bottleneck of the target model, which can exhaust memory in parallel scaling scenarios. As shown in Figure 1, as the number of sampling increases, the memory overhead of the target model tends to grow due to KV cache accumulation. This effect becomes more pronounced in target models when attempting to scale the number of requests from the draft model. During real-world deployment on the SGLang server (Zheng et al., 2024), high-concurrency sampling can lead to memory peaks that easily exceed the GPU’s maximum capacity, potentially causing the server to crash. Therefore, it is crucial to constrain the request budget from the draft model to the target model within a manageable range.
Figure 2: Comparison of naive and asynchronous speculative decoding.
The second bottleneck is the need for the target model to rank all samples from each draft model’s inference step before deciding which steps to accept. This presents a synchronization time bottleneck. As illustrated in Fig. 2, during the multi-turn sampling process (with a sampling quantity of 3 at each step), if the target model aims to reject the paths with the lowest two confidences, a ranking operation must be performed at each step. Given the limited computational and memory resources, the target model needs to prioritize processing the most important requests. Especially in test-time scaling, when combining sequential and parallel scaling, the synchronization overhead from precise budget control and the pursuit of globally optimal ranking is amplified. Although this issue has been widely discussed in the context of tool calls (Gim et al., 2024; Ginart et al., 2024), it has not been formally proposed in the context of test-time scaling.
To analyze the synchronization bottleneck and address the two challenges mentioned above, we first introduce a novel variant of arithmetic intensity called asynchronous arithmetic intensity to analyze the system bottleneck, and then explore conformal prediction (Vovk et al., 2005; 2003; Romano et al., 2020; Lei et al., 2018) for prediction ranking to design the asynchronous algorithm. In our formulation, conformal prediction defines a prediction set $C_{\alpha}$; sampling in $C_{\alpha}$ are rejected, while sampling outside are accepted. This yields a distribution-free guarantee that the right sampling is retained (i.e., lies outside $C_{\alpha}$) with high probability, enabling asynchronous test-time scaling. This paper presents the following contributions:
  • We propose asynchronous arithmetic intensity, a new metric designed to characterize and quantify throughput/latency bottlenecks that emerge in test-time scaling scenarios.
  • We introduce conformal prediction to address the memory bottleneck risks and the challenge of prediction ranking.
  • We propose A1 and implement a training-free, lossless acceleration, achieving a 56.7x speedup in test-time scaling and a 4.14x improvement in throughput through sequential and parallel test-time scaling.
2 Preliminary
In this section, we first introduce how to construct the prediction set in classic setup, and then present our setup.
Classic Setup.
Formally, let
$D_{\mathrm{cal}}=\bigl((X_{1},Y_{1}),\dots,(X_{n},Y_{n})\bigr)$ (1)
denote the calibration dataset. Each pair $(X_{i},Y_{i})$ for $i=1,\dots,n$ is a data point, consisting of an $X_{i}$ (the input question for the $i$-th example) and a ground-truth denoted as $Y_{i}$. The symbol $n$ represents the total number of calibration dataset.
$(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}).$ (2)
In adaptive conformal prediction (Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2023), the conformity score for each candidate path is computed using a softmax function:
$s_{i}^{k}=\frac{\exp(-\ell(X_{i},\hat{Y}_{i}^{k}))}{\sum_{j=1}^{m}\exp(-\ell(X_{i},\hat{Y}_{i}^{j}))},$ (3)
where -$\ell$ denotes the negative log-likelihood loss. Next, the global conformity threshold $\tau$ is obtained by computing the $p$-quantile over all candidate scores:
$\tau=Q_{p}\left(\left\{s_{i}^{j}\mid i=1,\dots,n;\ j=1,\dots,m\right\}\right)$ (4)
where
$p=\frac{\lceil(n+1)(1-\alpha)\rceil}{n},$ (5)
and $\alpha\in(0,1)$ is the user-specified miscoverage rate. The quantile $Q_{p}$ returns the $p$-th percentile smallest score, ensuring that the resulting prediction set includes the ground-truth with probability at least $1-\alpha$.
Problem Setup.
In the asynchronous test-time scaling setup, we leverage a draft model for fast sampling and delegate verification to a slower target model. Unlike classical rejection sampling (Chen et al., 2023), the target model evaluates the sampled chains of thought and proactively initiates its own resampling whenever a candidate is rejected, rather than passively awaiting new inputs. This rejection sampling process is repeated until the maximum number of scaling turns is reached, which inevitably introduces additional overhead in this setting. To mitigate synchronization and memory overhead on the target model’s server, we aim to predict and control the rejection rate by constructing prediction sets with bounded size. Conformal prediction offers a principled way to achieve this. Given a predefined $\alpha$, we estimate a confidence level such that the ground-truth $y$ falls within the prediction set $C_{\alpha}(Y)$ with probability at least $1-\alpha$:
$\vskip-5.69054pt\mathbb{P}(y\in C_{\alpha}(Y))\geq 1-\alpha,$ (6)
where $\alpha$ is conventionally interpreted as the significance level (e.g., 0.05 corresponding to 95% confidence). In this work, however, we reinterpret $\alpha$ as the rejection rate of the target model.
Ordinal Classification.
However, employing conformal prediction also suffers from the effect of synchronization (Zhu et al., 2025). Specifically, in conformal prediction, the conformal scores assigned to the measured objects are typically required to be normalized, as shown in Eq.3, and this normalization process inherently introduces a bottleneck to parallelization. Moreover, in typical inference engines (Zheng et al., 2024; Kwon et al., 2023), particularly those with asynchronous scheduling, obtaining the normalized scores for all sampling in different batches can be challenging. To address these challenges, we reformulate the task of constructing prediction sets as a ordinal classification (Dey et al., 2023; Xu et al., 2023), where the size of the prediction set corresponds to a reject rate of $\alpha$. Formally, we aim to ensure:
$\mathbb{P}(\tilde{y}^{i}\in C_{\alpha}(Y))\geq 1-\alpha,\quad\forall i\in\{1,\dots,n\times m\},$ (7)
where $\mathbb{P}(\tilde{y}^{i}\in C_{\alpha}(Y))$ denotes the probability that the $i$-th candidate step $\tilde{y}^{i}$ lies within the prediction set $C_{\alpha}$, and $m$ represents the number of sampled steps. This procedure provides marginal coverage, meaning that the coverage guarantee holds on average over the distribution of test inputs. The stronger notion of conditional coverage aims to ensure
$\mathbb{P}(\tilde{y}^{i}\in C_{\alpha}(Y)\mid X=x)\geq 1-\alpha,\quad\forall i\in\{1,\dots,m\},\ \forall x.$ (8)
that is, to provide a probabilistic guarantee for the sampled outputs corresponding to each input instance. Our setup focuses on developing asynchronous algorithms for ordinal prediction, with the construction of the set ensuring that its size exactly matches the predefined size under both marginal and conditional coverage.
(a) Sampling latency.
(b) End-to-end latency per turn.
Figure 3: Execution cost comparison between synchronous and asynchronous test-time scaling.
3 Challenge and Design
To identify performance bottlenecks in classic setup, we focus on the arithmetic intensity (Spector & Re, 2023), which quantifies the utilization of arithmetic units. Arithmetic intensity is defined as:
$I=\frac{F}{B},$ (9)
where $F$ denotes the number of floating-point operations (FLOPs), and $B$ denotes the number of bytes accessed.
3.1 Q1: What are the Emerging Performance Bottlenecks?
(a) Arithmetic intensity vs. sampling size.
(b) Asynchronous arithmetic intensity $r$ comparison.
Figure 4: Analysis of arithmetic intensity.
Speculative decoding (Leviathan et al., 2023) accelerates inference by overlapping computation with speculative memory accesses, enabling multiple draft tokens to be validated in parallel. Its main bottleneck is parallel score computation (Yin et al., 2024), making the process computation-bound.
Building upon this perspective, asynchronous scaling can be seen as an even more aggressive parallelization strategy. The target model validates far more tokens in parallel than in speculative decoding, which intensifies the prefill bottleneck and results in total computation time far exceeding total memory access time, as illustrated by the comparison between the green and yellow lines in Fig. 4(a). In addition, synchronization delays caused by varying sampling speeds across draft steps directly degrade computational resource utilization, which is not captured by Eq. 9. As shown in Fig. 3(b), synchronization overhead grows exponentially with the number of sampling turns. In the parallel scaling setting (Fig. 3(a)), this overhead increases linearly with the number of concurrent paths.
To this end, we observe Fig. 4(a) that increasing the sampling size naturally raises arithmetic intensity (with memory access time being negligible). However, due to synchronization overhead, the overall inference time of the system is increasing. To account for synchronization costs within arithmetic intensity, we define an asynchronous arithmetic intensity $r$:
$r=\frac{T_{c}}{T_{m}+T_{s}}=\frac{t_{c}\times F}{t_{m}\times B+T_{s}}\approx\frac{T_{c}}{T_{s}},$ (10)
where $T_{c}$ is computation time, $T_{m}$ is memory access time, $t_{c}$ and $t_{m}$ are the per-unit costs of computation and memory access, respectively. It can be observed from Fig. 4(b) that under synchronization settings, $r$ decreases as the sampling size increases which indicates that, in test-time scaling supported by speculative decoding, synchronization overhead emerges as the primary bottleneck.
3.2 Q2: How is the prediction set constructed?
Online Calibration.
Adaptive conformal prediction typically relies on a held-out calibration set to determine the threshold $\tau$, which ensures coverage guarantees. However, in the test-time scaling setup, held-out examples are generally unavailable during inference, making it impractical to reserve a dedicated calibration split. To address this limitation, we propose an online calibration strategy. Specifically, $m$ outputs are pre-sampled for each input in the test set, yielding $(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m})$. We then relax the strict requirement that conformity scores sum to one, and instead directly define:
$s_{i}^{k}=-\ell(X_{i},\hat{Y}_{i}^{k}).$ (11)
This formulation is used to estimate marginal conformal p-values for rejection sampling:
$p_{\xi}^{k}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbf{1}(s_{\xi}^{k}\leq s_{i}^{j})+1}{nm+1}$ (12)
In this formulation, $s_{\xi}^{k}$ denotes the conformity score of the test-time candidate $\hat{Y}_{\xi}^{k}$, which represents the $k$-th sample of the $\xi$-th input on the test set, and $s_{i}^{j}$ are the scores from the calibration set $(X_{i},\hat{Y}_{i}^{j})$. The indicator function $\mathbf{1}(\cdot)$ returns 1 when the condition is satisfied. The resulting conformal $p$-value governs rejection sampling: a candidate is accepted if $p_{n+1}^{k}>\alpha$, ensuring that only high-confidence outputs are retained.
This procedure guarantees marginal coverage at level $1-\alpha$ by selecting the $\lceil p\cdot n\cdot m\rceil$-th smallest score. Conditional coverage can be achieved by adjusting the quantile threshold by selecting the $\lceil p\cdot m\rceil$-th smallest score. A detailed comparison of these two approaches is provided in Appendix A.4.
Budget Prediction.
Let $B$ denote the predefined budget (i.e., the number of candidates to reject). Given a test-time input $X_{i}$, we sample $m$ candidate CoTs $\big(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}\big)$. For simplicity, we will not distinguish between conditional coverage and marginal coverage in what follows, and mweishenmweishen represents the total number of samples; we then compute their corresponding $p$-values $p_{i}^{1},\dots,p_{i}^{m}$}.
Importantly, this sampling and evaluation process is conducted asynchronously: each candidate $\hat{Y}_{i}^{k}$ is generated independently and evaluated for its $p$-value without requiring synchronization with other candidates. As a result, the outputs implicitly exhibit a descending order:
$p_{i}^{1}\geq p_{i}^{2}\geq\cdots\geq p_{i}^{m}.$ (13)
The ordered set can be partitioned using a threshold to construct the prediction set, by directly comparing each candidate’s $p$-value with the miscoverage threshold $\alpha$. Specifically, the prediction set includes all candidates whose $p$-values satisfy:
$\mathcal{C}_{\alpha}(Y_{i})=\left\{\hat{Y}_{i}^{k}:k\in\{1,\ldots,m\},p_{i}^{k}>\alpha\right\}.$ (14)
This formulation ensures that the selected candidates meet the conformal coverage criterion and can be safely retained. Equivalently, this can be interpreted as rejecting the top-$B$ lowest-$p$ candidates.
Figure 5: Asynchronous test-time scaling pipeline. The green box illustrates parallel scaling and follows the rejection sampling procedure, while the blue box illustrates sequential scaling.
3.3 How to perform rejection sampling via conformal prediction?
We adopt a three-stage sampling pipeline, illustrated in Fig. 5, to realize rejection sampling with a target rejection rate $\alpha$.
Draft Model Sampling.
Given input tokens $x_{1:n-1}$, the draft model proposes $m$ candidate continuations of length $K_{d}$, denoted $\tilde{y}_{n:n+K_{d}-1}^{j}$, by sampling from the same draft distribution $q_{d}$:
$\tilde{y}_{n:n+K_{d}-1}^{j}\sim q_{d}(\cdot\mid x_{1:n-1}),\quad j=1,\dots,m,$ (15)
with the joint specified tokenwise as
$q_{d}\!\left(\tilde{y}_{n:n+K_{d}-1}^{j}\mid x_{1:n-1}\right)=\prod_{i=0}^{K_{d}-1}q_{d}\!\left(\tilde{y}_{n+i}^{j}\,\middle|\,x_{1:n-1},\tilde{y}_{n:n+i-1}^{j}\right),\quad j=1,\dots,m.$ (16)
Verification.
For each candidate sampling $\tilde{y}_{n:n+K_{d}-1}^{j}$, we score it under the target model $q_{t}$ by computing the logits/likelihood $q_{t}(\tilde{y}_{n:n+K_{d}-1}^{j}\mid x_{1:n-1})$ and converting to a conformity score. Using the calibration set, we compute a $p$-value for each candidate and reject those inside $\mathcal{C}_{\alpha}$; otherwise accept.
Target Model Sampling.
Although classical rejection sampling discards the rejected samples from the draft model and resamples entirely from the target model, to save the token budget we proceed as follows. In each turn, the per-round target-side token budget is $K_{t}$: for each accepted candidate in $\mathcal{C}_{\alpha}$, we let the target model $q_{t}$ continue generation using that candidate as a prefix for up to $K_{t}$ tokens, stopping earlier if an end token is encountered.
Termination.
We iterate the above stages until a final answer is detected, the maximum number of turns is reached, or the overall token limit is exceeded. As highlighted in Fig. 5, increasing the number of turns enables sequential test-time scaling (blue box), while increasing the number of candidates per round enables parallel test-time scaling (green box).
Table 1: The comparison when the draft and target models are from different families, under the setting on marginal coverage (Mar Cov.) and conditional coverage (Con Cov.). indicates that both are reasoning models, whereas indicates that only the target model is a reasoning model. The results highlighted in red indicate the lossless acceleration method.
Dataset Draft Model (DM) / Target Model (TM) Type Mar Cov. Con Cov. DM Baseline TM Baseline
MATH100
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 87.0 94.0 83.0 96.0
Qwen2.5-7B-Instruction / QwQ-32B 86.0 96.0 84.0 96.0
Llama-3.1-8B-Instruct / QwQ-32B 87.0 95.0 75.0 96.0
DeepSeek-R1-Distill-Llama-8B / QwQ-32B 96.0 96.0 85.0 96.0
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 86.0 95.0 83.0 96.0
Qwen2.5-7B-Instruction / s1.1-32B 89.0 96.0 84.0 96.0
Llama-3.1-8B-Instruct / s1.1-32B 79.0 85.0 75.0 96.0
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 88.0 95.0 85.0 96.0
AIME24
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 86.7 66.7 60.0 86.7
Qwen2.5-7B-Instruction / QwQ-32B 46.7 33.3 33.3 86.7
Llama-3.1-8B-Instruct / QwQ-32B 33.3 40.0 13.3 86.7
DeepSeek Llama-3.1-8B-Instruct / QwQ-32B 80.0 80.0 80.0 86.7
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 73.3 80.0 60.0 86.7
Qwen2.5-7B-Instruction / s1.1-32B 40.0 33.3 33.3 86.7
Llama-3.1-8B-Instruct / s1.1-32B 26.7 40.0 13.3 86.7
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 73.3 73.3 80.0 86.7
AIME25
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 53.3 46.7 40.0 73.3
Qwen2.5-7B-Instruction / QwQ-32B 33.3 40.0 26.7 73.3
Llama-3.1-8B-Instruct / QwQ-32B 26.7 33.3 20.0 73.3
DeepSeek Llama-3.1-8B-Instruct / QwQ-32B 66.7 60.0 46.7 73.3
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 60.0 60.0 40.0 66.7
Qwen2.5-7B-Instruction / s1.1-32B 33.3 40.0 26.7 66.7
Llama-3.1-8B-Instruct / s1.1-32B 26.7 20.0 20.0 66.7
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 66.7 60.0 46.7 66.7
AMC23
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 88.0 90.0 74.0 94.0
Qwen2.5-7B-Instruction / QwQ-32B 76.0 72.0 68.0 94.0
Llama-3.1-8B-Instruct / QwQ-32B 68.0 68.0 44.0 94.0
DeepSeek Llama-3.1-8B-Instruct / QwQ-32B 92.0 94.0 80.0 94.0
DeepSeek-R1-Distill-Qwen-1.5B / s1.1-32B 86.0 86.0 74.0 96.0
Qwen2.5-7B-Instruction / s1.1-32B 74.0 78.0 68.0 96.0
Llama-3.1-8B-Instruct / s1.1-32B 54.0 64.0 44.0 96.0
DeepSeek Llama-3.1-8B-Instruct / s1.1-32B 92.0 93.0 80.0 96.0
4 Experiment
4.1 Asynchronous Test-time scaling across different model families.
We provide the hyperparameters used in our experiments in Appendix 4. Table 1 reports the results of asynchronous test-time scaling when the draft model (DM) and target model (TM) come from different families. We evaluate performance under two distinct settings: marginal coverage (Mar Acc.), which measures whether the predefined budget and prediction set align on average across test inputs, and conditional coverage (Con Acc.), which imposes a stricter requirement that the guarantee holds for each individual input instance. We have the following key takeaways: i) Speculative decoding can match the performance of the target model itself. This approach effectively reduces computational overhead while maintaining high-quality outputs, offering a feasible solution for scaling inference. ii) The most challenging datasets, AIME24/25, show strong performance in marginal coverage setup, while the other two datasets (MATH100 and AMC23) demonstrate superior results in conditional coverage setup. This highlights that while marginal coverage allocates more computational resources to the most difficult parts of the tasks effectively, conditional coverage ensures more reliable results at the individual input level, especially in simpler tasks, ensuring that each question is answered correctly. iii) This shows that even though reasoning models typically consume more tokens during inference, when constrained to the same token budget, using a reasoning model as the draft model yields stronger scaling performance compared to a non-reasoning model.
Table 2: The comparison when the draft and target models are from the same families.
Dataset
Draft Model / Target Model
Mar Acc.
Con Acc.
SpecReason
SpecThink
MATH100 Qwen2.5-7B/32B-Instruct 86.0 81.0 73.7 79.0
s1.1-7B/32B 88.0 87.0 73.7 85.0
DeepSeek-R1-Distill-Qwen-1.5B/32B 88.0 87.0 76.7 84.0
Skywork-OR1-7B/32B 88.0 89.0 75.8 70.0
AIME24 Qwen2.5-7B/32B-Instruct 33.3 40.0 33.3 33.3
s1.1-7B/32B 66.7 73.3 26.7 40.0
DeepSeek-R1-Distill-Qwen-1.5B/32B 86.7 80.0 66.7 66.7
Skywork-OR1-7B/32B 86.7 80.0 73.3 60.0
AIME25 Qwen2.5-7B/32B-Instruct 40.0 33.3 40.0 26.7
s1.1-7B/32B 53.3 53.3 40.0 33.3
DeepSeek-R1-Distill-Qwen-1.5B/32B 60.0 53.3 35.7 46.7
Skywork-OR1-7B/32B 60.0 53.3 53.3 40.0
AMC23 Qwen2.5-7B/32B-Instruct 72.0 70.0 72.0 74.0
s1.1-7B/32B 82.0 78.0 78.0 76.0
DeepSeek-R1-Distill-Qwen-1.5B/32B 92.0 88.0 80.0 82.0
Skywork-OR1-7B/32B 96.0 94.0 86.0 82.0
(a) Left: Conditional Coverage.
(b) Right: Marginal Coverage.
Figure 6: The budget prediction accuracy shows the accuracy of budget prediction for sampling lengths of 500 and 700 with a budget of 500 tokens during the online calibration phase under 64-sample and 16-sample settings. The experiment is conducted with a rejection rate $\alpha=0.25$.
4.2 Performance of Budget prediction
In this section, we present the accuracy of budget prediction under the settings of marginal coverage and conditional coverage. It demonstrates whether our method can accurately control the number of target model interventions in rejection sampling. This directly shows the accuracy of conformal prediction in estimating the rejection rate.
Marginal Coverage.
In Figure 6(b), we report the accuracy of the target-model intervention rate under marginal coverage, where the rejection rate is predicted at the dataset level. Budget-prediction accuracy across the full dataset is high—especially with the 64-sample configuration, whose absolute error stays within $5\%$. With $K_{d}=500$ tokens for calibration and $K_{d}=500$ for sampling, the error remains within $2\%$. This directly highlights the importance of constructing a diverse calibration set for maintaining high conformal-prediction accuracy.
Conditional Coverage.
In practice, we require precise per-batch budget control—rather than a single aggregate budget over the entire test set. Because the target-model server has limited capacity and cannot process all requests concurrently, inference is executed in batches, and the token budget must be enforced for each batch to satisfy the load constraint. In Figure 6(a), we report the accuracy of the target model intervention rate under conditional coverage. Under online calibration with a rejection rate of $25\%$, when $K_{d}=500$ in both the calibration and sampling stages, the 16-sample and 64-sample settings achieve similar accuracy. By contrast, when the calibration stage uses $K_{d}=500$ but the sampling stage uses $K_{d}=700$, the 64-sample setting attains significantly higher budget-prediction accuracy. This indicates that increasing the number of parallel samples can improve budget-prediction accuracy when the sampling token budget differs from the calibration token budget (i.e., under a calibration–sampling token-budget mismatch).
4.3 Asynchronous Test-time Scaling within The Same Model Families
In this section, we examine the performance across models within the same family, including both reasoning and non-reasoning variants in Table 2. In this setting, since the target model and draft models share the same vocabulary, we can compare against baselines that are only applicable to models within the same family, such as Speculative Thinking (Yang et al., 2025), denoted as SpecThink. Our findings are as follows: i): When the draft and target models belong to the same family, larger target models consistently outperform their smaller counterparts, confirming the benefit of intra-family consistency. ii): Across datasets, DeepSeek and Skywork show the strongest gains on challenging benchmarks (AIME, AMC), while Qwen2.5 performs competitively on MATH100 but lags significantly on harder tasks. iii): Moreover, s1.1 achieves moderate improvements, usually surpassing Qwen but not reaching the level of Skywork or DeepSeek. iv): Finally, speculative baselines (SpecReason and SpecThink) generally underperform compared with target model, suggesting that their effectiveness remains limited in reasoning-intensive settings.
Figure 7: Token budget at a maximum length of 20000 and 16 samples.
4.4 Analysis of token budget and latency
(a) Latency increases with the number of samples.
(b) Throughput variation under the 16-sample setting.
Figure 8: Analysis of latency and throughput trade-offs under different sampling and token budget.
Latency and Throughput.
In this part, we analyze the trade-offs between latency and throughput under different sampling and token budget settings. i): As shown in Figure 8(a), the latency of SpecReason and SpecThinking consistently increases with the number of samples, highlighting the cost of scaling up sampling. In contrast, our method significantly reduces the sampling latency and achieves the lowest inference latency under the condition coverage setting. ii): The overhead of online calibration is nearly negligible, particularly at larger sample sizes. iii): Meanwhile, Figure 8(b) illustrates how throughput varies with the per-sample token budget for methods such as SpecReason and SpecThinking, revealing diminishing returns as the token budget becomes large. Our method is able to maintain high throughput even under very large budgets, especially in the marginal coverage setting. Overall, these results (Figure 8) provide insights into the balance between efficiency and performance when designing inference strategies. Under the setting of 16 samples and a maximum token budget of 8192 per sample, we achieved a 56.7x speedup in inference and a 4.14x throughput improvement compared to the baseline.
Token Consumption.
We present in Figure 7 the token consumption under the 16-sample setting. This includes sampling performed solely by the target model or the draft model; as well as asynchronous sampling under both condition coverage and marginal coverage settings. Compared with the two baselines of the draft model and the target model, our method can significantly reduce token consumption, especially under the condition coverage setting, since it enables budget prediction at the instance level.
5 conclusion
We presented A1 (Asynchronous test-time scaling with conformal prediction), a framework that addresses the core inefficiencies of test-time scaling in LLMs. By refining arithmetic intensity and introducing online calibration with a rejection sampling pipeline, A1 effectively controls rejection rates while reducing latency and memory overhead. Experiments on multiple reasoning benchmarks confirm that A1 achieves better efficiency and reliability than speculative baselines. This work establishes A1 as a practical and principled approach for scalable LLM inference, with potential extensions to dynamic adaptation and real-world deployment.
Appendix A Appendix
A.1 Related Work
Test-time Scaling.
Recent works explore test-time scaling—the idea that increasing computation during inference can be more effective than scaling model size (Snell et al., 2024; Wu et al., 2024). A common strategy is sequential scaling, adopted in models like OpenAI o1 (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025). Other approaches (Muennighoff et al., 2025; Yan et al., 2025) use supervised fine-tuning to match a fixed compute budget. In parallel, parallel scaling (Chen et al., 2025; Zeng et al., 2025; Pan et al., 2025a) improves throughput by distributing inference across replicas or devices, offering latency gains but introducing challenges in memory overhead.
Speculative decoding
Speculative decoding(Li et al., 2024a; Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023) is an emerging technique for accelerating LLM inference, which is traditionally limited by slow, sequential autoregressive sampling and memory bandwidth constraints. There are three main strategies for sampling draft tokens: token-level sampling(Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023), where the large model directly verifies the token outputs of the draft model; feature-level sampling (Cai et al., 2024; Li et al., 2024a; b), which verifies generation paths using intermediate representations; and step-level sampling (Pan et al., 2025b; Yang et al., 2025), which operates at a coarser granularity by validating multiple tokens or computation steps together to improve throughput.
Asynchronous Tool Calling.
The synchronization issue in batch inference with tool calls (Zhu et al., 2025) is a known obstacle to efficient reasoning. However, it remains underexplored in the context of speculative decoding—particularly when large model inference is treated as a form of tool call. In asynchronous scheduling, controlling the frequency of large model intervention is challenging due to synchronization overhead. Recent approaches (Ginart et al., 2024; Gonzalez-Pumariega et al., 2025) employ event-driven finite-state machine architectures to manage asynchronous tool calls more flexibly and efficiently.
Conformal Prediction.
To avoid synchronization and to accurately predict the request budget in the scaling process, we introduce conformal prediction (Derhacobian et al., ; Angelopoulos et al., 2020; Huang et al., 2023) to provide a theoretical guarantee for the budget of times our target model intervenes. The prediction set is then used to ensure that the large model’s interventions remain consistent with the desired coverage and reliability, aligning with the validation process. However, these methods all require the model to perform a complete softmax operation (which requires synchronization), and this becomes challenging in modern inference engines with asynchronous scheduling mechanisms, thus conflicting with these methods. Some online conformal prediction algorithms (Areces et al., ; Bhatnagar et al., 2023) attempt to ensure the coverage of future data in the context of online learning.
A.2 Experimental Setup
We evaluate a diverse set of draft models, including DeepSeek-R1-Distill-Qwen-1.5B (Guo et al., 2025), DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025), Qwen2.5-7B-Instruct (Team, 2024), Llama-3.1-8B (Dubey et al., 2024), s1.1-7B (Muennighoff et al., 2025), and Skywork-OR1-7B (He et al., 2025). Each draft model is paired with one of the large target models: QwQ-32B (Team, 2025), s1.1-32B (Muennighoff et al., 2025), Qwen2.5-32B-Instruct (Team, 2024), DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025), or Skywork-OR1-32B (He et al., 2025). We use SpecReason (Pan et al., 2025b) and Speculative Thinking (Yang et al., 2025) as baselines for speculative decoding, with a maximum length of 8192. Both of them focus on acceleration under serial scaling.
Our evaluation covers four reasoning benchmarks. We use 100 randomly sampled problems from MATH (Hendrycks et al., 2021) (denoted as MATH100) for grade-school arithmetic word problems, AIME24 (Art of Problem Solving, ) and AIME25 (OpenCompass, ) for high-school competition-level mathematics, and the first 50 problems from AMC23 (math-ai, ) for the American Mathematics Competitions. These datasets require multi-step reasoning and are particularly suitable for testing the effectiveness of asynchronous sampling with rejection. To ensure consistency, we set a token budget of 8192 across all settings and adopt deterministic decoding with temperature set to zero.
Similar to prior work (Yue et al., 2025), the best@16 metric we calculate is intended to measure the upper bound of performance for both the method and the baseline model. Unless otherwise specified, the miscoverage parameter is set to $\alpha=0.4$, ensuring that the prediction sets are constructed with statistical guarantees. We use SGLang (Zheng et al., 2024) version 0.4.3.post4 as the inference engine. The sampling temperature is set to 0.8. We set the target model’s per-turn token budget to $K_{t}=500$ and the draft model’s per-turn token budget to $K_{d}=500$.
A.3 Controlling the rejection rate in multi-turn interactions
(a) PPL variation with increasing turns.
(b) Take-over counts across multiple turns.
Figure 9: Analysis of model behavior in multi-turn interactions. The left subfigure shows how the Perplexity (PPL) of draft and target models evolves as the number of turns increases, while the right subfigure presents the take-over counts across turns.
In this section, we study the variation of model perplexity with sequential scaling and the change in the number of take-overs by the large model during rejection sampling. This directly reflects whether the conformal prediction algorithm we adopt can control the intervention rate of the target model within an acceptable range. i): According to Figure 9(a), each generation by the draft model leads to an increase in the overall PPL, while the intervention of the target model effectively mitigates this trend. As the number of turns increases, the PPL of the scaling process can be gradually reduced, which indirectly results in a decrease in the rejection rate. ii): Meanwhile, in Figure 9(b), we directly visualize the change in the number of take-overs by the target model as the number of interaction turns increases. We observe that with sequential scaling, the target model overall maintains a rejection rate around a fixed level, which gradually decreases and eventually approaches zero. iii): The target model continuously adjusts the convergence behavior of the draft model’s perplexity during sequential scaling.
A.4 Definition and Proof
Proposition 1.
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable random variables from the test dataset, and $\xi\sim\text{Uniform}\{1,2,\ldots,n\}$ represents randomly sampling one data point from the test dataset, where $k$ denotes the $k$-th sample of that data point, then the marginal conformal $p$-values defined as,
$p_{\xi}^{k}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{m}\mathbf{1}(s_{\xi}^{k}\leq s_{i}^{j})+1}{nm+1}$ (17)
is valid in the sense that for the miscoverage rate $\alpha\in(0,1)$, we have
$\mathbb{P}(p_{\xi}^{k}\leq\alpha)\leq\alpha.$ (18)
Moreover, if the conformity scores $\{s_{i}^{j}\}_{i=1,j=1}^{n,m}$ are distinct surely, we have,
$p_{\xi}^{k}\sim U\left\{\frac{1}{nm+1},\frac{2}{nm+1},\ldots,1\right\}.$ (19)
Proof of Proposition 1.
Suppose, for any given values of conformity scores, $v_{1},\ldots,v_{nm+1}$, they can be rearranged as $\tilde{v}_{1}<\cdots<\tilde{v}_{\ell}$ with repetitions $n_{i}$ of $\tilde{v}_{i}$ such that $\sum_{i=1}^{\ell}n_{i}=nm+1$. Let $E_{v}$ denote the event of $\{s_{1}^{1},s_{1}^{2},\ldots,s_{n}^{m},s_{\xi}^{k}\}=\{v_{1},\ldots,v_{nm+1}\}$.
Then, under $E_{v}$, for $i=1,\ldots,\ell$, we have
$\mathbb{P}(s_{\xi}^{k}=\tilde{v}_{i}\mid E_{v})=\frac{n_{i}}{nm+1},$ (20)
due to the exchangeability of conformity scores.
We also note that under $E_{v}$ and $s_{\xi}^{k}=\tilde{v}_{i}$ we have from Equation equation 17,
$p_{\xi}^{k}=\frac{\sum_{l=1}^{i}n_{l}}{nm+1}.$ (21)
Then, for any $\alpha\in[0,1]$ and $i=1,\ldots,\ell$, we have
$\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v},s_{\xi}^{k}=\tilde{v}_{i})=\begin{cases}0&\text{if }\alpha<\frac{\sum_{l=1}^{i}n_{l}}{nm+1},\\ 1&\text{otherwise}.\end{cases}$ (22)
Thus, for any $i=1,\ldots,\ell$ and $\frac{\sum_{l=1}^{i-1}n_{l}}{nm+1}\leq\alpha<\frac{\sum_{l=1}^{i}n_{l}}{nm+1}$, we have
$\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v})$ $=\sum_{l=1}^{\ell}\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v},s_{\xi}^{k}=\tilde{v}_{l})\cdot\mathbb{P}(s_{\xi}^{k}=\tilde{v}_{l}\mid E_{v})$ (23)
$=\frac{\sum_{l=1}^{i-1}n_{l}}{nm+1}\leq\alpha.$ (24)
By taking the expectation over the above inequality, it follows that the conformal $p$-value $p_{\xi}^{k}$ is marginally valid.
Specifically, if conformity scores $\{s_{i}^{j}\}_{i=1,j=1}^{n,m}\cup\{s_{\xi}^{k}\}$ are distinct surely, then $\ell=nm+1$ and $n_{i}=1$ for $i=1,\ldots,nm+1$. Thus,
$\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v})=\frac{i-1}{nm+1},\quad\text{if }\frac{i-1}{nm+1}\leq\alpha<\frac{i}{nm+1},$ (25)
that is, $p_{\xi}^{k}\mid E_{v}\sim U\left\{\frac{1}{nm+1},\frac{2}{nm+1},\ldots,1\right\}$. This completes the proof. Next, we address the theoretical part that guarantees conditional coverage. ∎
Proposition 2.
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable random variables from the test dataset, then for any sample $y\in\mathcal{Y}$, given a conditioning set $\mathcal{I}_{y}\subseteq\{0,\ldots,m-1\}$ for each $X_{i}$ constructed based on the specific sample $y$ for a test input $x$ and $Y_{\xi}=y$, the corresponding conditional conformal $p$-value as defined in Equation equation 17, is conditionally valid in the sense that for any $\alpha\in[0,1]$,
$\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid\mathcal{I}_{y},Y_{\xi}=y)\leq\alpha.$ (26)
Moreover, if $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, we have that conditional on $\mathcal{I}_{y}$ and $Y_{\xi}=y$,
$p_{\xi}^{k}\sim U\left\{\frac{1}{m+1},\frac{2}{m+1},\ldots,1\right\},$ (27)
where $m=|\mathcal{I}_{y}|$ is the size of the conditioning set for output $y$.
Proof.
For any given sample $y\in\mathcal{Y}$, the corresponding conditional conformal $p$-value is given by
$p_{\xi}^{k}=\frac{1}{m+1}\left(\sum_{i=1}^{n}\sum_{j\in\mathcal{I}_{y}}\mathbf{1}\{s_{i}^{j}\leq s_{\xi}^{k}\}+1\right),$ (28)
where $\mathcal{I}_{y}\subseteq\{0,\ldots,m-1\}$ is the conditioning set for sample $y$, $m=|\mathcal{I}_{y}|$, $s_{i}^{j}$ represents the conformity score for the $j$-th candidate of the $i$-th test instance, and $s_{\xi}^{k}$ is the conformity score for the $k$-th candidate of the new test instance.
Given $\mathcal{I}_{y}$ and $Y_{\xi}=y$, the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}\cup\{s_{\xi}^{k}\}$ are exchangeably distributed, which follows from the assumption that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeably distributed and the construction of candidate outputs.
Using similar arguments as in the proof of Proposition 1, for any given values of conformity scores $v_{1},\ldots,v_{nm+1}$, suppose that they can be arranged as $\tilde{v}_{1}<\cdots<\tilde{v}_{\ell}$ with repetitions $m_{i}$ of $\tilde{v}_{i}$ such that $\sum_{i=1}^{\ell}m_{i}=nm+1$. Let $E_{v}$ denote the event $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}\cup\{s_{\xi}^{k}\}=\{v_{1},\ldots,v_{nm+1}\}$.
Then, given $E_{v}$, $\mathcal{I}_{y}$, and $Y_{\xi}=y$, we have
$\mathbb{P}(s_{\xi}^{k}=\tilde{v}_{i}\mid E_{v},\mathcal{I}_{y},Y_{\xi}=y)=\frac{m_{i}}{nm+1}$ (29)
for $i=1,\ldots,\ell$, due to exchangeability of the conformity scores.
Note that given $E_{v}$, $\mathcal{I}_{y}$, $Y_{\xi}=y$, and $s_{\xi}^{k}=\tilde{v}_{i}$, we have from Equation equation 28,
$p_{\xi}^{k}=\frac{\sum_{j=1}^{i}m_{j}}{nm+1}.$ (30)
Thus, for any $\alpha\in[0,1]$ and $i=1,\ldots,\ell$,
$\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v},\mathcal{I}_{y},Y_{\xi}=y,s_{\xi}^{k}=\tilde{v}_{i})=\begin{cases}0&\text{if }\alpha<\frac{\sum_{j=1}^{i}m_{j}}{nm+1},\\ 1&\text{otherwise}.\end{cases}$ (31)
Then, for any given $i=1,\ldots,\ell$ and $\frac{\sum_{j=1}^{i-1}m_{j}}{nm+1}\leq\alpha<\frac{\sum_{j=1}^{i}m_{j}}{nm+1}$, we have
$\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v},\mathcal{I}_{y},Y_{\xi}=y)$ (32)
$=\sum_{j=1}^{\ell}\mathbb{P}(p_{\xi}^{k}\leq\alpha\mid E_{v},\mathcal{I}_{y},Y_{\xi}=y,s_{\xi}^{k}=\tilde{v}_{j})\cdot\mathbb{P}(s_{\xi}^{k}=\tilde{v}_{j}\mid E_{v},\mathcal{I}_{y},Y_{\xi}=y)$ (33)
$=\frac{\sum_{j=1}^{i-1}m_{j}}{nm+1}\leq\alpha.$ (34)
By taking expectation, it follows that $p_{\xi}^{k}$ is conditionally valid given $Y_{\xi}=y$. This completes the proof. ∎
Discussion.
After proving the validity of individual conformal p-values, in order to obtain the rejection rate for the entire test set, that is, to ensure that the overall error rate is controlled when simultaneously testing K hypotheses, we propose the following Proposition.
Proposition 3.
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable random variables, and let $\xi\sim\text{Uniform}\{1,2,\ldots,n\}$ represent a randomly selected test instance, then A1 based on marginal conformal $p$-values all provide simultaneous coverage guarantees across the entire test dataset at level $1-\alpha$, i.e.,
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))\geq 1-\alpha.$ (35)
Specifically, if the conformity scores $\{s_{i}^{j}\}_{i=1,j=1}^{n,m}$ are distinct surely, then for A1, we also have,
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))\geq 1-\alpha+\frac{1}{(n-1)m+1}.$ (36)
Proof of Proposition 3.
Consider A1 based on marginal conformal $p$-values. Note that among the tested hypotheses $H_{1},\ldots,H_{m}$, there is exactly one hypothesis $H_{Y_{\xi}}$ to be true. Thus, the probability that all samples are correctly covered by A1 satisfies:
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))$ $=\mathbb{P}(\text{accept }H_{Y_{\xi}})$ (37)
$=1-\mathbb{P}(\text{reject }H_{Y_{\xi}})$ (38)
$\geq 1-\mathbb{P}(p_{\xi}^{Y_{\xi}}\leq\alpha)$ (39)
$\geq 1-\alpha,$ (40)
where the last inequality follows by Proposition 1.
Specifically, for A1, if the conformity scores $\{s_{i}^{j}\}_{i=1,j=1}^{n,m}$ are distinct surely, by Proposition 1, we have
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))$ $=\mathbb{P}(\text{accept }H_{Y_{\xi}})$ (41)
$=1-\mathbb{P}(p_{\xi}^{Y_{\xi}}\leq\alpha)$ (42)
$\geq 1-\left(\alpha-\frac{1}{(n-1)m+1}\right)$ (43)
$=1-\alpha+\frac{1}{(n-1)m+1},$ (44)
which gives the desired result. ∎
Theorem 1.
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable random variables, and let $\xi\sim\text{Uniform}\{1,2,\ldots,n\}$ represent a randomly selected test instance, then the prediction set $\mathcal{C}(X_{\xi})=\left\{\hat{Y}_{\xi}^{k}\mid p_{\xi}^{k}>\alpha\right\}$ determined by A1 both satisfy
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))\geq 1-\alpha.$ (45)
Proof.
Note that the prediction set is given by $\mathcal{C}(X_{\xi})=A_{1}\cap A_{2}$. Thus, by Proposition 1,
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))\geq\mathbb{P}(p_{\xi}^{Y_{\xi}}>\alpha)\geq 1-\alpha.$ (46)
Similarly, its prediction set is given by
$\mathcal{C}(X_{\xi})=\{y\in\mathcal{Y}:p_{\xi}^{y}>\alpha\}.$ (47)
By Proposition 1, it is easy to check that
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))=\mathbb{P}(p_{\xi}^{Y_{\xi}}>\alpha)\geq 1-\alpha.$ (48)
Specifically, if the conformity scores $\{s_{i}^{j}\}_{i=1,j=1}^{n,m}$ are distinct surely, we have
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))=1-\mathbb{P}(p_{\xi}^{Y_{\xi}}\leq\alpha)\leq 1-\alpha+\frac{1}{(n-1)m+1}.$ (49)
This completes the proof. ∎
Discussion.
Now that we have completed the proof of marginal coverage, we proceed to prove the conditional coverage for the entire test dataset. Under the exchangeability assumption, we have:
Proposition 4.
Under the same exchangeability assumption as in Proposition 2, A1 based on conditional conformal $p$-values $p_{\xi}^{k}$ all provide conditional coverage guarantees for the entire test dataset at level $1-\alpha$, i.e., for any $y\in\mathcal{Y}$,
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid Y_{\xi}=y)\geq 1-\alpha.$ (50)
Specifically, if the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, then for A1 based on $p_{\xi}^{k}$, we have that for any $y\in\mathcal{Y}$ and $\mathcal{I}_{y}\subseteq\{0,\ldots,m-1\}$,
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid Y_{\xi}=y,\mathcal{I}_{y})\geq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}.$ (51)
Proof.
Consider Procedure 1-3 based on conditional conformal $p$-values. For any $y\in\mathcal{Y}$, given $Y_{\xi}=y$, the conditional coverage probability for the entire test dataset satisfies:
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid Y_{\xi}=y)$ $=1-\mathbb{P}(\text{reject }H_{y}\mid Y_{\xi}=y)$ (52)
$\geq 1-\mathbb{P}(p_{\xi}^{y}\leq\alpha\mid Y_{\xi}=y)$ (53)
$\geq 1-\alpha,$ (54)
where the inequalities follow from the definitions of Procedure 1-3 and Proposition 2.
Specifically, for Procedure 3, if the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, then by Proposition 2, the coverage probability conditional on $\mathcal{I}_{y}$ and $Y_{\xi}=y$ satisfies:
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid\mathcal{I}_{y},Y_{\xi}=y)$ $=1-\mathbb{P}(\text{reject }H_{y}\mid\mathcal{I}_{y},Y_{\xi}=y)$ (55)
$=1-\mathbb{P}(p_{\xi}^{y}\leq\alpha\mid\mathcal{I}_{y},Y_{\xi}=y)$ (56)
$\geq 1-\left(\alpha-\frac{1}{(n-1)|\mathcal{I}_{y}|+1}\right)$ (57)
$=1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}.$ (58)
This completes the proof. ∎
Theorem 2.
Under the same exchangeability assumption as in Theorem 1, the prediction set $\mathcal{C}(X_{\xi}\mid y)$ based on conditional conformal $p$-values $p_{\xi}^{k}$ satisfies
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid Y_{\xi}=y)\geq 1-\alpha$ (59)
for any $y\in\mathcal{Y}$. Specifically, for the prediction set $\mathcal{C}(X_{\xi}\mid y)$ based on $p_{\xi}^{k}$, if the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, we have
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid Y_{\xi}=y,\mathcal{I}_{y})\leq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}$ (60)
for any $y\in\mathcal{Y}$ and $\mathcal{I}_{y}\subseteq\{0,\ldots,m-1\}$.
Proof.
By using Proposition 4 and similar arguments as in the proof of Theorem 1, the prediction sets $\mathcal{C}(X_{\xi}\mid y)$ based on the conditional conformal $p$-values $p_{\xi}^{k}$ all satisfy:
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid\mathcal{I}_{y},Y_{\xi}=y)\geq 1-\alpha$ (61)
for any $y\in\mathcal{Y}$.
Specifically, if the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, we have:
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid\mathcal{I}_{y},Y_{\xi}=y)$ $=\mathbb{P}(p_{\xi}^{y}>\alpha\mid\mathcal{I}_{y},Y_{\xi}=y)$ (62)
$\leq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1},$ (63)
which gives the desired result. ∎